OpenStudy (anonymous):

$200 in the account and has earned a total of $8 in interest in two years, what is the rate of the account?

OpenStudy (texaschic101):

I = PRT is the original formula, but since we are looking for R, we can change it to : I/PT = R I = 8 P = 200 T = 2 now we sub I /PT = R 8/(200)(2) = R 8/400 = R .02 = R -- change to percent 2% = R

OpenStudy (anonymous):

thanks so much !! (:

OpenStudy (texaschic101):

sure thing :)